OpenStudy (anonymous):

please help!! 1medal and fan Use the Multiplication Property of Equality to complete the following statement: If 4x / 2=4 then _____ 4x=8 4x * 2=8 4=4x * 2 4=4x / 2

Directrix (directrix):

(4x / 2) = 4 Multiply both sides of the equation by 2. What do you get? @knowel

Directrix (directrix):

What is 2/1 times (4x)/2?

OpenStudy (anonymous):

i got 2 and 4x

OpenStudy (anonymous):

for(4x / 2) = 4 i got 2 and for 2/1 times (4x)/2 i got 4x

Directrix (directrix):

2/1 times (4x)/2 =